# The OATMEAL Survey. II. The 3D spin-orbit obliquity of an eccentric transiting brown dwarf in the Ruprecht 147 open cluster

### v1

@@ Abstract
We present new analysis of the CWW 89 system as part of the Orbital Architectures of Transiting Massive Exoplanets And Low-mass stars (OATMEAL) survey. The CWW 89 system is a member of the 2.8 Gyr old Ruprecht 147 (NGC 6774) cluster and features two stars, CWW 89A (EPIC 219388192) and CWW 89B, with the primary hosting a transiting brown dwarf. We use in-transit, highly precise radial velocity measurements with the Keck Planet Finder (KPF) to characterize the Rossiter-McLaughlin (RM) effect and measure the projected spin-orbit obliquity $|\lambda|=1.4\pm 2.5^{\circ}$ and the full 3D spin-orbit obliquity of the brown dwarf to be $\psi=15.1^{+15.0^{\circ}}_{-10.9}$. This value of $\lambda$ implies that the brown dwarf’s orbit is prograde and well-aligned with the equator of the host star, continuing the trend of transiting brown dwarfs showing a preference for alignment ($\lambda\approx 0^{\circ}$) regardless of the stellar effective temperature. We find that this contrast with the transiting giant planet population, whose spin-orbit alignments depend on host $T_{\rm eff}$, shows an increasingly clear distinction in the formation and orbital migration mechanisms between transiting giant planets and transiting brown dwarfs like CWW 89Ab. For this system in particular, we find it plausible that the brown dwarf may have undergone coplanar high-eccentricity migration influence by CWW 89B.
@@ 1 Introduction
Our understanding of the tidal realignment of close-in planetary orbits around main sequence stars has substantially improved since the first measurements of spin-orbit obliquities were made for gas giant planets (D. Queloz et al., 2000a; J. N. Winn et al., 2005; A. H. M. J. Triaud et al., 2010; T. Hirano et al., 2011; S. Albrecht et al., 2012). Here, we define “spin-orbit obliquity” to mean the angle between the rotation axis of the primary star and the orbital axis of the companion(s). With recent works by J. Zak et al. (2024); J. I. Espinoza-Retamal et al. (2024); E. Knudstrup et al. (2024); X.-Y. Wang et al. (2024); J. Zhang et al. (2025), the number of obliquity measurements made for giant planets now numbers in the dozens, with a wide range of aligned to misaligned systems. A relatively coherent story is emerging for the orbital evolution of these systems and it is centered on a general theory of the interior structures of stars that host giant planets. As main sequence stars that are more massive than the Sun evolve, expand in radius, and decrease in effective temperature, their outer envelopes transition from being predominantly radiative to convective. R. P. Kraft (1967) first established a temperature threshold at which this change occurs, now known as the Kraft Break ($T_{\rm eff}=6250$K). From this, a trend in the projected spin-orbit alignments, or obliquities $\lambda$, for giant planets around the Kraft Break has emerged: Stars hotter than the Kraft Break show a population of planets with no preferential value of $\lambda$ whereas stars cooler than the Kraft Break yield a companion planet population with preferentially aligned obliquities ($\lambda=0^{\circ}$) (e.g. J. N. Winn et al., 2010; M. Rice et al., 2022; J. Rusznak et al., 2024). Recent observations of subgiant stars that have crossed the Kraft break during their evolution confirmed that alignments of giant planets are strongly tied to presence of convective envelopes (e.g. N. Saunders et al., 2024).
@@
The more massive siblings to giant planets, brown dwarfs (BDs), have recently seen a boom in their known transiting population in large part thanks to the TESS mission (G. R. Ricker et al., 2015) and efforts from ground-based facilities like the Next Generation Transit Survey (NGTS) (P. J. Wheatley et al., 2018). Recent works from (B. A. Henderson et al., 2024a; N. Vowell et al., 2025; A. Larsen et al., 2025, and references therein) show more than 50 known transiting brown dwarfs and like with their giant planet siblings, astronomers have pursued measurements of their spin-orbit obliquities for insight into their orbital histories. However, in contrast to the dozens spin-orbit alignments measured for transiting giant planets, only 9 such measurements exist for transiting BDs. Every measurement of $\lambda$ for transiting BDs thus far indicates a preference for an aligned orbit (A. H. M. J. Triaud et al., 2009; R. J. Siverd et al., 2012; A. H. M. J. Triaud et al., 2013; G. Zhou et al., 2019; S. Giacalone et al., 2024; T. Ferreira dos Santos et al., 2024; M. Brady et al., 2025; L. Doyle et al., 2025; C. M. Johns-Krull et al., 2008).
@@
One widely-used technique to measure $\lambda$ for transiting system is the Rossiter-McLaughlin (RM) Effect (D. Queloz et al., 2000b; A. H. M. J. Triaud, 2018). This requires taking spectra of the host star during a transit to observe the radial velocity (RV) anomaly induced by the transiting companion blocking varying amounts of red- or blue-shifted light depending on the orbital configuration relative to the orientation of the host star’s spin axis. The RM effect alone can only offer an estimate of the projected spin-orbit alignment $\lambda$, but in some cases, the system supplies all of the ingredients needed to determine the full 3D obliquity $\psi$, namely both the orbital inclination of the transiting companion and the stellar spin axis angle, which is determined from combining the projected rotational velocity of the star with its rotation period (often observed via photometric modulations in its light curve).
@@
Here we present an RM measurement of $\lambda$ for the transiting BD CWW 89Ab. CWW 89Ab’s environment distinguishes it from other BD systems with measured obliquities. The CWW 89 system receives its name from J. L. Curtis et al. (2013a), who catalog it and over 100 other stars as members of the Ruprecht 147 open cluster. This makes CWW 89Ab one of four known transiting BDs that reside in a stellar association or cluster (the other three being AD 3116b in Praesepe, RIK 72b in Upper Scorpius, and HIP 330609b in MELANGE 6 from E. Gillen et al. (2017), T. J. David et al. (2019), and N. Vowell et al. (2023) respectively), and the oldest BD to reside in such an environment at $2.76\pm 0.61$ Gyr (G. Torres et al., 2018). CWW 89 consists of a primary G-dwarf star, CWW 89A, and a secondary M-dwarf, CWW 89B, orbiting at a 24.9 AU (81 mas) projected separation (T. G. Beatty et al., 2018). A handful of transiting BD systems comprise 3 or more objects, typically a primary star, a transiting BD companion, and a more distant outer stellar companion, but CWW 89 is the only multi-star transiting BD system featuring a G-dwarf star as its primary.
@@
The paper proceeds as follows: Section 2 details past and new RV measurements used in the orbital solution and characterization of the RM effect, respectively. Section 3 details our RM model, calculation of the stellar spin axis angle, and the determination of the full 3D obliquity. Section 4 outlines considerations for the emerging trend in spin-orbit alignments for the transiting brown dwarf population.
@@ 2.1 Keck Planet Finder in-transit RVs
From 07:10 UT to 12:19 UT on 2024 July 30, we observed a transit of the CWW 89Ab using the Keck Planet Finder (KPF; S. R. Gibson et al., 2024a) with the Keck I telescope at the W. M. Keck Observatory located on Maunakea, HI, USA. To ensure sufficient pre- and post-transit baseline, we observed the star continuously from approximately 1 hour before ingress to 1 hour after egress. We took a series of spectra with KPF at 600 s exposure times each, resulting in 20 in-transit observations at a SNR of 60-90 across the 500-600nm wavelength range. The goal was to maintain a SNR at this level or higher while keeping good coverage of the transit so that the RV anomaly induced by the RM effect would be well-characterized. The RVs are calculated using the KPF data reduction pipeline (S. R. Gibson et al., 2024b). The complete series of KPF RVs are given in Table 1. We note that KPF produces RVs in a green and red CCD, taking the RV values from different orders in the echelle spectrum.
Based on reported the stellar parameters in G. Nowak et al. (2017) and T. W. Carmichael et al. (2019) of $M_{\star}=1.1\rm\,M_{\sun}$, $R_{\star}=1.0\rm\,R_{\sun}$, $T_{\rm eff}=5755$K, $\log{g}=4.5$, and $\rm[Fe/H]=0.2$, we assume CWW 89A is a sunlike star. We only use the RVs derived from the green CCD as these are at a higher SNR than those of the red CCD for this star by a factor of $\lesssim 2.5$.
@@ 2.2 Literature RVs from TRES and FIES
Past works by G. Nowak et al. (2017) and T. W. Carmichael et al. (2019) have published orbital solutions (i.e. out-of-transit spectra) for CWW 89A. These data are used to remove the Keplerian motion from the in-transit RVs newly collected with KPF. These data aid in the detailed analysis of the RM effect to account for the out-of-transit motion of the brown dwarf.
T. W. Carmichael et al. (2019) provide 18 RVs measured from spectra taken with the Tillinghast Reflector Echelle Spectrograph (TRES) located at the Fred Lawrence Whipple Observatory at Mt. Hopkins, AZ, USA. These data were acquired from 2015 September 24 to 2016 May 30. Here, we simply use the values published in Table 2 of T. W. Carmichael et al. (2019).
@@ 2.3 K2 Data
We use photometric time series from the K2 missions as part of our joint modeling of the RM effect. The K2 data are acquired from the Mikulski Archive for Space Telescopes and are processed following the methods outlined in T. W. Carmichael et al. (2019). We note that TESS is scheduled to observe CWW 89A in its sector 92 (May 2025) of Extended Mission 2, so those data are not include here.
@@ 3 Analysis
We jointly model the KPF, TRES, and FIES RVs, K2 light curves, and host star broadband SED photometry using a modified version [^1] Modified EXOFASTv2 of EXOFASTv2 (J. D. Eastman et al., 2019) where instead of using a simplified approximation for the RM effect from, Y. Ohta et al. (2005), we employ a more robust calculation from T. Hirano et al. (2011), which accounts for several subtle effects, including stellar rotation, macroturbulence, thermal broadening, pressure broadening, and instrumental broadening.
@@
Figure 1: Top: RV orbital solution for CWW 89A using TRES (T. W. Carmichael et al., 2019) and FIES (G. Nowak et al., 2017) data. Bottom: KPF in-transit RV data showing the RV anomaly induced by the Rossiter-McLaughlin Effect.
@@
We set either uniform $\mathcal{U}[a,b]$ or Gaussian $\mathcal{G}[a,b]$ priors on our input parameters (see Table 3). These priors have mean values $a$ with width values $b$. We use spectroscopic priors on $T_{\rm eff}$ and [Fe/H] from T. W. Carmichael et al. (2019), zero-point corrected (L. Lindegren et al., 2018) parallax priors from Gaia DR3, and an upper limit on V-band extinction ($A_{V}$, E. F. Schlafly & D. P. Finkbeiner, 2011). J. L. Curtis et al. (2013b) confirm the CWW 89 system to be a probable member of the Ruprecht 147 open cluster, so we additionally set the cluster’s age of $2.76\pm 0.61$ (G. Torres et al., 2018) as a prior. We let the RV offsets $\gamma$ for each instrument be free parameters and we include an RV jitter term, $\sigma_{j}$, for each RV dataset to account for the surface activity of the star. The RV jitter term is based on a white noise model implemented in EXOFASTv2. We use values from T. W. Carmichael et al. (2019) for the following parameters to set as starting points for our fit: the orbital period, mid-transit time, orbital inclination, and the companion-to-host radius ratio.
@@
For the RM model, we model the limb-darkening for KPF assuming a $V$ bandpass wavelength range. We fix $v_{\alpha}$ to 0 (ignore differential rotation) and we use scaling relations for $v_{\zeta}$ (macroturbulence) from A. P. Doyle et al. (2014) and for $v_{\xi}$ (microturbulence) from H. Bruntt et al. (2010). Both $v_{\zeta}$ and $v_{\xi}$ have uniform priors. From this joint analysis of the stellar properties, transit photometry, and RVs, we find a projected obliquity angle of $|\lambda|=1.4\pm 2.5^{\circ}$ for CWW 89Ab.
@@
All stellar parameters we derive are consistent within $<2\sigma$ of those determined in G. Nowak et al. (2017) and T. W. Carmichael et al. (2019). Our full list of derived parameters are shown in Table 3. We highlight our derived $T_{\rm eff}=5680\pm 86$K for CWW 89A as this is especially relevant in our discussion of the realignment timescale for CWW 89Ab. A summary of other basic parameters we derive: $M_{\star}=1.02\pm 0.03\rm\,M_{\sun}$, $R_{\star}=1.07\pm 0.02\rm\,R_{\sun}$, [Fe/H]=0.07$\pm$0.09, $M_{b}=37.3\pm 1.6\rm\,M_{\rm J}$, $R_{b}=0.96\pm 0.04\rm\,R_{\rm J}$, $P_{\rm orb}=5.29$ days.
@@
Table 2: Table of transiting brown dwarf projected obliquities $\lambda$ and 3D obliquities $\psi$. Values for CWW 89Ab are from this work.
@@
HATS-70b $12.9\pm 1.8$ $8.9\pm 5.6$ 1
GPX-1b $19.7\pm 1.6$ $6.9\pm 10.0$ 2
CoRoT-3b $22.3\pm 1.0$ $37.6^{+10.0}_{-22.3}$ 3
KELT-1b $27.3\pm 0.5$ $2.0\pm 16$ 4
CWW 89Ab $39.2\pm 1.1$ $1.4\pm 2.5$ $15.1^{+15.0}_{-10.9}$
WASP-30b $62.5\pm 1.2$ $7.0^{+19.0}_{-27.0}$ 5
@@
TOI-2533b $74.9\pm 5.3$ $7.0\pm 14.0$ 8
footnotetext: References: 1) G. Zhou et al. (2019), 2) S. Giacalone et al. (2024), 3) A. H. M. J. Triaud et al. (2009), 4) R. J. Siverd et al. (2012), 5) A. H. M. J. Triaud et al. (2013), 6) L. Doyle et al. (2025), 7) M. Brady et al. (2025), 8) T. Ferreira dos Santos et al. (2024)
Figure 2: The posterior distribution for the stellar inclination angle $i_{\star}$ for CWW 89A. We use the $P_{\rm rot}$ measured from G. Nowak et al. (2017) and the sampling method first presented in K. Masuda & J. N. Winn (2020) and described in B. P. Bowler et al. (2023) to derive an $i_{\star}=74.7^{+9.4^{\circ}}_{-18.6}$. We interpret this as aligned with the orbital inclination $i_{0}$.

### v2

@@ Abstract
We present new analysis of the CWW 89 system as part of the Orbital Architectures of Transiting Massive Exoplanets And Low-mass stars (OATMEAL) survey. The CWW 89 system is a member of the 2.8 Gyr old Ruprecht 147 (NGC 6774) cluster and consists of two stars, CWW 89A (EPIC 219388192) and CWW 89B, with the primary hosting a transiting brown dwarf, CWW 89Ab. We use in-transit, highly precise radial velocity measurements with the Keck Planet Finder (KPF) to characterize the Rossiter-McLaughlin (RM) effect and measure the projected spin-orbit obliquity $|\lambda|=1.4\pm 2.5^{\circ}$ and the full 3D spin-orbit obliquity of the brown dwarf to be $\psi=15.1^{+15.0^{\circ}}_{-10.9}$. This value of $\lambda$ implies that the brown dwarf’s orbit is prograde and well-aligned with the equator of the host star, continuing the trend of transiting brown dwarfs showing a preference for spin-orbit alignment ($\lambda\approx 0^{\circ}$) regardless of the stellar effective temperature. This contrast with the transiting giant planet population, whose spin-orbit alignments depend on host $T_{\rm eff}$, showing an increasingly clear distinction in the formation and orbital migration mechanisms between transiting giant planets and transiting brown dwarfs like CWW 89Ab. For this system in particular, we find it plausible that the brown dwarf may have undergone coplanar high-eccentricity migration influenced by CWW 89B.
@@ 1 Introduction
Our understanding of the tidal realignment of close-in planetary orbits around main sequence stars has substantially improved since the first measurements of spin-orbit obliquities were made for gas giant planets (first_rm; winn2005; triaud2010; hirano2011; albrecht2012). Here, we define “spin-orbit obliquity” to mean the angle between the rotation axis of the primary star and the orbital axis of the companion. With recent works by zak2024; espinoza2024; knudstrup2024; wang2024; jingwen2025, the number of obliquity measurements made for giant planets now numbers in the dozens, with a wide range of aligned to misaligned systems. A relatively coherent story is emerging for the orbital evolution of these systems and it is centered on a general theory of the interior structures of stars that host giant planets. As main sequence stars that are more massive than the Sun evolve, expand in radius, and decrease in effective temperature, their outer envelopes transition from being predominantly radiative to convective. kraft1967 first established a temperature threshold at which this change occurs, now known as the Kraft Break at a $T_{\rm eff}=6250$K. From this, a trend in the projected spin-orbit alignments, or obliquities $\lambda$, for giant planets around the Kraft Break has emerged: Stars hotter than the Kraft Break show a population of planets with no preferential value of $\lambda$ whereas stars cooler than the Kraft Break yield a companion planet population with preferentially aligned obliquities ($\lambda=0^{\circ}$) (e.g. winn2010; rice2022; espinoza2023_CHEM; rusznak2024). Recent observations of subgiant stars that have crossed the Kraft break during their evolution confirmed that alignments of giant planets are strongly tied to presence of convective envelopes (e.g. saunders24).
@@
The more massive siblings to giant planets, brown dwarfs (BDs), have recently seen a boom in their known transiting population in large part thanks to the TESS mission (tess) and efforts from ground-based facilities like the Next Generation Transit Survey (NGTS) (ngts). Recent works from (henderson2024; vowell2025; larsen2025, and references therein) show more than 50 known transiting brown dwarfs and like with their giant planet siblings, astronomers have pursued measurements of their spin-orbit obliquities for insight into their orbital histories. However, in contrast to the dozens spin-orbit alignments measured for transiting giant planets, only 9 such measurements exist for transiting BDs. Every measurement of $\lambda$ for transiting BDs thus far indicates a preference for an aligned orbit (xo3b_1; corot3_RM; kelt1b; wasp30_rm; hats70b; oatmeal1; dosSantos2024; brady2025; doyle2025).
@@
One widely-used technique to measure $\lambda$ for transiting system is the Rossiter-McLaughlin (RM) Effect (first_rm; rm_effect). This requires taking spectra of the host star during a transit to observe the radial velocity (RV) anomaly induced by the transiting companion blocking varying amounts of red- or blue-shifted light depending on the orbital configuration relative to the orientation of the host star’s spin axis. The RM effect alone only offers an estimate of the projected spin-orbit alignment $\lambda$, but in some cases, the system supplies all of the ingredients needed to determine the full 3D obliquity $\psi$, namely the orbital inclination $i_{0}$ of the transiting companion and the stellar spin axis angle $i_{\star}$, which is determined from combining the projected rotational velocity of the star with its rotation period (often observed via photometric modulations in its light curve).
@@
Here we present a Rossiter-McLaughlin measurement of $\lambda$ for the transiting BD CWW 89Ab (also known as EPIC 219388192b and TOI-7497). CWW 89Ab’s environment distinguishes it from other BD systems with measured obliquities. The CWW 89 system receives its name from curtis2013, who catalog it and over 100 other stars as members of the Ruprecht 147 open cluster. This makes CWW 89Ab one of four known transiting BDs that reside in a stellar association or cluster. The other three are AD 3116b in Praesepe, RIK 72b in Upper Scorpius, and HIP 330609b in MELANGE 6 from ad3116, david19_bd, and vowell2023, respectively. CWW 89Ab is the oldest BD to reside in such an environment at $2.76\pm 0.61$ Gyr (r147_age). CWW 89 consists of a primary G-dwarf star, CWW 89A, and a secondary M-dwarf, CWW 89B, orbiting at a 24.9 AU (81 mas) projected separation (cww89a). A handful of transiting BD systems comprise 3 or more objects, typically a primary star, a transiting BD companion, and a more distant outer stellar companion, but CWW 89 is the only multi-star transiting BD system featuring a G-dwarf star as its primary.
@@
The paper proceeds as follows: Section 2 details past and new RV measurements used in the orbital solution and characterization of the RM effect. Section 3 details our RM model, calculation of the stellar spin axis angle, and the determination of the full 3D obliquity. Section 4 outlines considerations for the emerging trend in spin-orbit alignments for the transiting brown dwarf population. Section 5 recapitulates our most important findings.
@@ 2.1 Keck Planet Finder in-transit RVs
From 07:10 UT to 12:19 UT on 2024 July 30, we observed a transit of the CWW 89Ab using the Keck Planet Finder (KPF; Gibson2024) with the Keck I telescope at the W. M. Keck Observatory located on Maunakea, HI, USA. To ensure sufficient pre- and post-transit baseline, we observed the star continuously from approximately 1 hour before ingress to 1 hour after egress. We took a series of spectra with KPF at 600 s exposure times each, resulting in 20 in-transit observations at a SNR of 60-90 across the 445-600nm wavelength range. The goal was to maintain a SNR at this level or higher while keeping good temporal sampling of the transit so that the RV anomaly induced by the RM effect would be well-detected. The RVs are calculated using the KPF data reduction pipeline (KPF_DRP). The complete series of KPF RVs are given in Table 1. We note that KPF produces RVs in a green and red CCD, taking the RV values from different orders in the echelle spectrum with the green CCD spanning 445-600nm and the red CCD spanning 600-870nm. We only use the RVs derived from the green CCD as these are at a higher SNR than those of the red CCD for this star by a factor of $\lesssim 2.5$.
Based on reported the stellar parameters in nowak17 and carmichael19 of $M_{\star}=1.1\rm\,M_{\sun}$, $R_{\star}=1.0\rm\,R_{\sun}$, $T_{\rm eff}=5755$K, $\log{g}=4.5$, and $\rm[Fe/H]=+0.2$, we assume CWW 89A is a sunlike star.
@@ 2.2 Literature RVs from TRES and FIES
Past works by nowak17 and carmichael19 have published orbital solutions (i.e. out-of-transit spectra) for CWW 89A. These data are used to remove the Keplerian motion from the in-transit RVs newly collected with KPF. These data aid in the detailed analysis of the RM effect to account for the out-of-transit motion of the brown dwarf.
carmichael19 provide 18 RVs measured from spectra taken with the Tillinghast Reflector Echelle Spectrograph (TRES) located at the Fred Lawrence Whipple Observatory at Mt. Hopkins, AZ, USA. These data were acquired from 2015 September 24 to 2016 May 30. Here, we simply use the values published in Table 2 of carmichael19.
@@ 2.3 K2 Data
We use photometric time series from the K2 missions as part of our joint modeling of the RM effect. The K2 data are acquired from the Mikulski Archive for Space Telescopes and are processed following the methods outlined in carmichael19.
@@ 3 Analysis
[…]
Figure 1: SED photometry of CWW 89A. Blue points are the model values from MIST (mist1; mist2; mist3). Red points are the input data from Gaia (gaia_dr3_magnitude_corr), 2MASS (2MASS), and WISE (WISE).
Table 2: Stellar magnitudes used in the SED fitting for CWW 89A. We impose a floor of 0.02 magnitudes on the uncertainty for the Gaia bandpasses to ensure that the SED model is not overconstrained by potentially underestimated photometric errors. footnotetext: References: 1 - gaia_dr3_magnitude_corr, 2 - 2MASS, 3 - WISE
Name Description Value Source
$G$. Gaia green. $12.36\pm 0.02$ 1
$G_{BP}$. Gaia blue. $12.76\pm 0.02$ 1
$G_{RP}$. Gaia red. $11.80\pm 0.02$ 1
$J$. 2MASS $J$. $11.07\pm 0.02$ 2
$H$. 2MASS $H$. $10.73\pm 0.02$ 2
$K_{S}$. 2MASS $K_{S}$. $10.67\pm 0.02$ 2
[…]
@@
Figure 2: Top: RV orbital solution for CWW 89A using TRES (carmichael19) and FIES (nowak17) data. Middle: KPF in-transit RV data showing the RV anomaly induced by the Rossiter-McLaughlin Effect. The y-axis units in the top and middle panels are in $\rm m\,s^{-1}$. Bottom: Phase-fold K2 transit photometry of CWW 89A with our EXOFASTv2 transit model overlaid.
@@
We set either uniform $\mathcal{U}[a,b]$ or Gaussian $\mathcal{G}[a,b]$ priors on our input parameters. The Gaussian priors have mean values $a$ with width values $b$ while the uniform priors use $a$ and $b$ and lower and upper bounds, respectively. We use spectroscopic priors on $T_{\rm eff}$ and [Fe/H] from carmichael19, zero-point corrected (Lindegren2018) parallax priors from Gaia DR3, and an upper limit on V-band extinction ($A_{V}$, av_priors). curtis2013 confirm the CWW 89 system to be a probable member of the Ruprecht 147 open cluster, so we additionally set the cluster’s age of $2.76\pm 0.61$ (r147_age) as a prior. We let the RV offsets $\gamma$ for each instrument be free parameters and we include an RV jitter term, $\sigma_{j}$, for each RV dataset to account for the surface activity of the star. The RV jitter term is based on a white noise model implemented in EXOFASTv2. We use values from carmichael19 for the following parameters to set as starting points for our fit: the orbital period, mid-transit time, orbital inclination, and the companion-to-host radius ratio.
@@
For the RM model, we model the limb-darkening for KPF assuming a $V$ bandpass wavelength range. We fix $v_{\alpha}$ to 0 (ignore differential rotation) and we use scaling relations for $v_{\zeta}$ (macroturbulence) from doyle2014_exofast and for $v_{\xi}$ (microturbulence) from bruntt2010_exofast. Both $v_{\zeta}$ and $v_{\xi}$ have uniform priors. From this joint analysis of the stellar properties, transit photometry, and RVs, we find a projected obliquity angle of $|\lambda|=1.4\pm 2.5^{\circ}$ for CWW 89Ab. We contextualize this measurement with others for transiting BD systems in Table 3.
@@
All stellar parameters we derive are consistent within $2\sigma$ of those determined in nowak17 and carmichael19. Our full list of derived parameters are shown in Table 4. We highlight our derived $T_{\rm eff}=5680\pm 86$K for CWW 89A as this is especially relevant in our discussion of the realignment timescale for CWW 89Ab. A summary of other system parameters we derive: $M_{\star}=1.02\pm 0.03\,M_{\sun}$, $R_{\star}=1.07\pm 0.02\,R_{\sun}$, $\rm[Fe/H]=+0.07\pm 0.09$, $M_{b}=37.3\pm 1.6\,M_{\rm J}$, $R_{b}=0.96\pm 0.04\,R_{\rm J}$, $P_{\rm orb}=5.29$ days.
@@
Table 3: Table of transiting brown dwarf projected obliquities $\lambda$ and 3D obliquities $\psi$. Values for CWW 89Ab are from this work. footnotetext: References: 1) hats70b, 2) oatmeal1, 3) corot3_RM, 4) kelt1b, 5) wasp30_rm, 6) doyle2025, 7) brady2025, 8) dosSantos2024
@@
HATS-70b $12.9\pm 1.8$ $8.9\pm 5.6$ - 1
GPX-1b $19.7\pm 1.6$ $6.9\pm 10.0$ - 2
CoRoT-3b $22.3\pm 1.0$ $37.6^{+10.0}_{-22.3}$ - 3
KELT-1b $27.3\pm 0.5$ $2.0\pm 16$ - 4
CWW 89Ab $39.2\pm 1.1$ $1.4\pm 2.5$ $15.1^{+15.0}_{-10.9}$ -
WASP-30b $62.5\pm 1.2$ $7.0^{+19.0}_{-27.0}$ - 5
@@
TOI-2533b $74.9\pm 5.3$ $7.0\pm 14.0$ - 8
Figure 3: The posterior distribution for the stellar inclination angle $i_{\star}$ for CWW 89A. We use the $P_{\rm rot}$ measured from nowak17 and the sampling method first presented in masuda2020 and described in bowler2023_inclination to derive an $i_{\star}=74.4^{+9.7^{\circ}}_{-18.3}$. We interpret this as aligned with the orbital inclination $i_{0}$.